Shelton Company has a debt−equity ratio of 1.28. Return on assets is 7.53 percent, and total equity is $640,000. What is the equity multiplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Equity multiplier times What is the return on equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Return on equity % What is the net income? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Net income $
Answer #1

debt−equity ratio=debt/equity

Hence debt=(1.28*640,000)=$819200

Total assets=debt+equity

=(819200+640,000)=$1,459,200

Equity multiplier=Total assets/equity

=(1,459,200/640,000)=2.28

ROA=Net income/Total assets

Net income=$1,459,200*7.53%

=$109,878(Approx).

ROE=Net income/equity

=$109,878/640,000

=17.17%(Approx).